### v1

@@ 4.1 Experiment Details
Datasets. ELE-40K comprises 44,094 image-mask pairs annotated with real-world transmission equipment and hazard scenarios, with 80% of the data allocated for training and 20% for validation. HQSeg-44K [21] is a comprehensive dataset combining six existing image segmentation datasets, including the training sets of DIS [34], ThinObject-5K [22], FSS [25], ECSSD [43], MSRA-10K [4], and DUT-OMRON [53]. The DIS, ThinObject-5K, COIFT, and HR-SOD datasets are used as validation sets. This results in a total of 44,320 annotated image-mask pairs. For the division of the dataset, we adopt the same configuration as HQ-SAM and PA-SAM. Additionally, we use COCO[42] for evaluating zero-shot segmentation.
Evaluation Metrics. Following HQ-SAM [21] for assessing high-quality segmentation, we adopt mean Intersection over Union (mIoU) and mean Boundary Intersection over Union (mBIoU) as metrics.
@@ 4.2 Results on ELE-40K
To comprehensively evaluate the effectiveness of ELE-SAM, we conduct comparative experiments on ELE-40K against various state-of-the-art segmentation models, including SAM [20], HQ-SAM [21], PA-SAM [50], U²Net [16], MvaNet [55], and IS-Net [34]. In addition to the overall performance, the detailed results on two primary categories, i.e., equipment and hazards, are also reported. The quantitative results in Tab. 2 demonstrate the superior mIoU and mBIoU performance of ELE-SAM. The results are analyzed in two main aspects: the impact of ELE-40K dataset and the performance superiority of ELE-SAM over other models.
Impact of ELE-40K Dataset. As demonstrated in Tab. 2, the performance of models like HQ-SAM [21], PA-SAM [50], and IS-Net-General-Use [34] when directly applied to ELE-40K without fine-tuning is significantly lower than their counterparts fine-tuned on ELE-40K (HQ-SAM-finetune [21], PA-SAM-finetune, and IS-Net). For instance, HQ-SAM achieves an mIoU of 44.8% on equipment and 70.7% on hazards, while its fine-tuned version obtains absolute improvements of 24.9% and 8.7%, respectively. Similarly, PA-SAM shows a substantial improvement from 36.3% to 71.3% mIoU on equipment after fine-tuning. This stark contrast indicates the distinctive challenge of the PTCHS task compare to general segmentation, thus highlighting the value of ELE-40K.
@@
Performance of ELE-SAM. ELE-SAM consistently outperforms SAM and its derivative models, as well as several state-of-the-art segmentation methods. As shown in Tab. 2, ELE-SAM outperforms the baseline SAM [20] and its fine-tuned version (SAM-finetune) [20], increasing the mIoU by 16.8% and mBIoU by 20.6% over SAM, and by 14.6% in mIoU and 17.6% in mBIoU over SAM-finetune in average. Furthermore, ELE-SAM achieves clear improvements over other prominent SAM-based models, particularly in equipment segmentation, with the mIoU improvement of 6.6% over HQ-SAM-finetune [21] and 5.0% over PA-SAM-finetune [50].
Compared to end-to-end segmentation models, ELE-SAM outperforms U²Net by 30.7% in mIoU and 30.2% in mBIoU, while surpassing IS-Net by 24.6% in mIoU and 24.1% in mBIoU. Similarly, ELE-SAM achieves a 29.9% and 30.0% improvement in mIoU and mBIoU over MvaNet. Overall, benefit from our designed CAPA and HFMD modules, which can preserve the object structure details in high-fidelity, ELE-SAM delivers promising performance on ELE-40K and significantly surpasses other competitors.
Visual Results. In addition to the quantitative results, we also provide some visualizations in Fig. 7a. Compared to the prominent HQ-SAM [21], ELE-SAM generates sharper and more complete segmentations with fewer false positives, particularly in challenging cases with fine structural details. The visualizations indicate that ELE-SAM excels at distinguishing power transmission equipment from complex background while preserving boundary integrity.
@@ 4.3 Effectiveness on High-Quality Segmentation
We further validate the effectiveness of our methods on high-quality segmentation for more general objects using HQSeg-44K [21], as shown in Tab. 3. Overall, ELE-SAM improves the average mIoU and mBIoU by 2.9% and 3.8% compared to the previous leading method. Regarding the four sub-sets, ELE-SAM obtains more significant enhancement on DIS and ThinObject. For instance, ELE-SAM outperforms PA-SAM [50] by 6.5% mIoU and 4.1% mBIoU on DIS, 2.6% mIoU and 6.0% mBIoU on ThinObject. Additionally, ELE-SAM surpasses HQ-SAM [21] by 9.4% mIoU and 7.6% mBIoU on DIS, 5.8% mIoU and 10.1% mBIoU on ThinObject. Since the two sub-sets contain abundant objects in mesh structure, such as steel cable bridge and iron fence, segmenting these objects in high quality requires larger mask feature resolution. HQ-SAM and PA-SAM only adopt the mask features with $256\times 256$ resolution, resulting in the perception loss of fine structures. As shown in Fig. 7b, HQ-SAM fails to segment the structure details of foreground objects. In comparison, ELE-SAM explores the generation and refinement of mask features in $512\times 512$ resolution, contributing to the substantial improvement of segmentation quality.
Figure 8: Comparison of visual results between SAM (top row) and ELE-SAM (bottom row) on the COCO[42] validation set under a zero-shot setting. FocalNet-DINO [57], trained on the COCO dataset is utilized as the box prompt generator. ELE-SAM demonstrates superior mask quality compared to SAM, achieving high-accuracy segmentation while maintaining robust zero-shot segmentation performance.
@@
Table 4: Comparison of zero-shot segmentation capabilities with SAM and its derivative models on the COCO [42] dataset. FocalNet-DINO [57] is used for generating box prompt.
@@
SAM [20] 48.5 75.5 52.7 63.9 53.1 34.1
SAM-Adapter [56] 44.8 69.5 48.1 63.9 47.8 29.0
@@ 4.4 Zero-Shot Comparison with SAM
To evaluate the generalization capability of our proposed ELE-SAM, we conduct zero-shot segmentation experiment on the COCO[42] dataset, comparing its performance against SAM and several derivative models. Following HQ-SAM [21], we use the model trained on HQSeg-44K for direct evaluation on COCO in the zero-shot manner. As summarized in Tab. 4, our method demonstrates superior performance across various metrics. Specifically, ELE-SAM achieves 50.6% Average Precision (AP), outperforming SAM (48.5%), HQ-SAM (49.5%), and PA-SAM [50] (49.9%). The results validate that our ELE-SAM also maintain the generalization capability. As illustrated in Fig. 8, ELE-SAM produces fewer artifacts and achieves more accurate segmentation, especially in complex background. It further underscore the robustness of ELE-SAM, making it a strong candidate for real-world applications which require high-quality segmentation without task-specific fine-tuning.
@@ 4.5 Ablation Study
In this section, we conduct ablation studies on ELE-40K and HQSeg-44K to investigate the effectiveness of High-Fidelity Mask Decoder (HFMD) and Context-Aware Prompt Adapter (CAPA). In HFMD, before the fusion process (the same as HQ-SAM [21]) between mask features and image fusion features, we plug in another light-weight refinement block (RB). The impacts of the additional refinement block and leveraging high-resolution mask features are discussed. Moreover, we showcase the effectiveness of CAPA over the original Prompt Adapter (PA).
@@
Impact of Different Backbones. When changing the size of ViT backbone, we find that ViT-L obtains the best performance on ELE-40K. The detailed metrics are listed in Tab. 7. While ViT-L outperforming ViT-B by a clear margin, the larger frozen ViT-H does not further promote the performance. It could be attributed to the over-smoothing issue [37] for deeper ViT.
@@ 5 Conclusion
[…]
Acknowledgements This work was supported in part by the National Natural Science Foundation of China under Grants U23B2048 and 62225113, in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017, in part by the Science and Technology Major Project of Hubei Province under Grant 2024BAB046, and in part by the Science and Technology Project of State Grid Corporation of China under Grant 5700-202458333A-2-1-ZX. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Declarations
  • Funding. Dr. Juhua Liu was supported by the National Natural Science Foundation of China under Grants U23B2048 and 62225113, in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017, in part by the Science and Technology Major Project of Hubei Province under Grant 2024BAB046, and in part by the Science and Technology Project of State Grid Corporation of China under Grant 5700-202458333A-2-1-ZX.
  • Competing Interests. The authors have no relevant financial or non-financial interests to disclose.
  • Data availability. All datasets used in this paper are publicly available. ELE-40K can be accessed at https://github.com/Hhaizee/ELE-SAM. HQSeg-44K [21] can be accessed at https://github.com/SysCV/SAM-HQ. COCO [42] can be accessed at https://github.com/cocodataset/cocoapi

### v2

@@ 4.1 Experiment Details
Datasets. ELE-40K comprises 44,094 image-mask pairs annotated with real-world transmission equipment and hazard scenarios, with 80% of the data allocated for training and 20% for validation. HQSeg-44K [sam_hq] is a comprehensive dataset combining six existing image segmentation datasets, including the training sets of DIS [qin2022highlyaccuratedichotomousimage], ThinObject-5K [9423134], FSS [Li_2020], ECSSD [7182346], MSRA-10K [6871397], and DUT-OMRON [6619251]. The DIS, ThinObject-5K, COIFT, and HR-SOD datasets are used as validation sets. This results in a total of 44,320 annotated image-mask pairs. For the division of the dataset, we adopt the same configuration as HQ-SAM and PA-SAM. Additionally, we use COCO[singh2024benchmarkingobjectdetectorscoco] for evaluating zero-shot segmentation.
Evaluation Metrics. Following HQ-SAM [sam_hq] for assessing high-quality segmentation, we adopt mean Intersection over Union (mIoU) and mean Boundary Intersection over Union (mBIoU) as metrics.
@@
Figure 8: Per-category mIoU comparison on ELE-40K. ELE-SAM achieves superior performance across 14 out of all 15 categories, with significant improvements on Equipment categories featuring fine structures. The bottom subplot shows instance counts.
@@ 4.2 Results on ELE-40K
To comprehensively evaluate the effectiveness of ELE-SAM, we conduct comparative experiments on ELE-40K against various state-of-the-art segmentation models, including SAM [kirillov2023segment], HQ-SAM [sam_hq], PA-SAM [xie2024pasampromptadaptersam], U²Net [fang2021denselynestedtopdownflows], MvaNet [yu2024multiview], and IS-Net [qin2022highlyaccuratedichotomousimage]. In addition to the overall performance, the detailed results on two primary categories, i.e., equipment and hazards, are also reported. The quantitative results in Tab. 2 demonstrate the superior mIoU and mBIoU performance of ELE-SAM. The results are analyzed in two main aspects: the impact of ELE-40K dataset and the performance superiority of ELE-SAM over other models.
Impact of ELE-40K Dataset. As demonstrated in Tab. 2, the performance of models like HQ-SAM [sam_hq], PA-SAM [xie2024pasampromptadaptersam], and IS-Net-General-Use [qin2022highlyaccuratedichotomousimage] when directly applied to ELE-40K without fine-tuning is significantly lower than their counterparts fine-tuned on ELE-40K (HQ-SAM-finetune [sam_hq], PA-SAM-finetune, and IS-Net). For instance, HQ-SAM achieves an mIoU of 44.8% on equipment and 70.7% on hazards, while its fine-tuned version obtains absolute improvements of 24.9% and 8.7%, respectively. Similarly, PA-SAM shows a substantial improvement from 36.3% to 71.3% mIoU on equipment after fine-tuning. This stark contrast indicates the distinctive challenge of the PTCHS task compare to general segmentation, thus highlighting the value of ELE-40K.
@@
Performance of ELE-SAM. ELE-SAM consistently outperforms SAM and its derivative models, as well as several state-of-the-art segmentation methods. As shown in Tab. 2, ELE-SAM outperforms the baseline SAM [kirillov2023segment] and its fine-tuned version (SAM-finetune) [kirillov2023segment], increasing the mIoU by 16.8% and mBIoU by 20.6% over SAM, and by 14.6% in mIoU and 17.6% in mBIoU over SAM-finetune in average. Furthermore, ELE-SAM achieves clear improvements over other prominent SAM-based models, particularly in equipment segmentation, with the mIoU improvement of 6.6% over HQ-SAM-finetune [sam_hq] and 5.0% over PA-SAM-finetune [xie2024pasampromptadaptersam]. Compared to end-to-end segmentation models, ELE-SAM outperforms U²Net by 30.7% in mIoU and 30.2% in mBIoU, while surpassing IS-Net by 24.6% in mIoU and 24.1% in mBIoU. Similarly, ELE-SAM achieves a 29.9% and 30.0% improvement in mIoU and mBIoU over MvaNet.
Per-Category Analysis. To provide a more fine-grained evaluation, we further examine the per-category mIoU performance in Fig. 8. ELE-SAM achieves particularly large improvements on challenging equipment categories. For example, for insulator and line, ELE-SAM surpasses PA-SAM-FineTune by 9.3% and 7.7% mIoU respectively, underscoring the enhanced ability to handle fine-structured objects.
Overall, benefiting from the proposed CAPA and HFMD modules, which effectively preserve object structural details with high fidelity, ELE-SAM achieves outstanding performance on the ELE-40K dataset and significantly outperforms existing competitors.
Visual Results. In addition to the quantitative results, we also provide some visualizations in Fig. 7(a). Compared to HQ-SAM [sam_hq], ELE-SAM generates sharper and more complete segmentations with fewer false positives, particularly in challenging cases with fine structural details. The visualizations indicate that ELE-SAM excels at distinguishing power transmission equipment from complex background while preserving boundary integrity.
@@ 4.3 Effectiveness on High-Quality Segmentation
We further validate the effectiveness of our methods on high-quality segmentation for more general objects using HQSeg-44K [sam_hq], as shown in Tab. 3. Overall, ELE-SAM improves the average mIoU and mBIoU by 2.9% and 3.8% compared to the previous leading method. Regarding the four sub-sets, ELE-SAM obtains more significant enhancement on DIS and ThinObject. For instance, ELE-SAM outperforms PA-SAM [xie2024pasampromptadaptersam] by 6.5% mIoU and 4.1% mBIoU on DIS, 2.6% mIoU and 6.0% mBIoU on ThinObject. Additionally, ELE-SAM surpasses HQ-SAM [sam_hq] by 9.4% mIoU and 7.6% mBIoU on DIS, 5.8% mIoU and 10.1% mBIoU on ThinObject. Since the two sub-sets contain abundant objects in mesh structure, such as steel cable bridge and iron fence, segmenting these objects in high quality requires larger mask feature resolution. HQ-SAM and PA-SAM only adopt the mask features with $256\times 256$ resolution, resulting in the perception loss of fine structures. As shown in Fig. 7(b), HQ-SAM fails to segment the structure details of foreground objects. In comparison, ELE-SAM explores the generation and refinement of mask features in $512\times 512$ resolution, contributing to the substantial improvement of segmentation quality.
Figure 9: Comparison of visual results between SAM (top row) and ELE-SAM (bottom row) on the COCO[singh2024benchmarkingobjectdetectorscoco] validation set under a zero-shot setting. FocalNet-DINO [zhang2022dinodetrimproveddenoising], trained on the $CO$ dataset is utilized as the box prompt generator. ELE-SAM demonstrates superior mask quality compared to SAM, achieving high-accuracy segmentation while maintaining robust zero-shot segmentation performance.
@@
Table 4: Comparison of zero-shot segmentation capabilities with SAM and its derivative models on the COCO [singh2024benchmarkingobjectdetectorscoco] dataset. FocalNet-DINO [zhang2022dinodetrimproveddenoising] is used for generating box prompt.
@@
SAM [kirillov2023segment] 48.5 75.5 52.7 63.9 53.1 34.1
SAM-Adapter [yang2023aimadaptingimagemodels] 44.8 69.5 48.1 63.9 47.8 29.0
@@ 4.4 Zero-Shot Comparison with SAM
To evaluate the generalization capability of our proposed ELE-SAM, we conduct zero-shot segmentation experiment on the COCO[singh2024benchmarkingobjectdetectorscoco] dataset, comparing its performance against SAM and several derivative models. Following HQ-SAM [sam_hq], we use the model trained on HQSeg-44K for direct evaluation on COCO in the zero-shot manner. As summarized in Tab. 4, our method demonstrates superior performance across various metrics. Specifically, ELE-SAM achieves 50.6% Average Precision (AP), outperforming SAM (48.5%), HQ-SAM (49.5%), and PA-SAM [xie2024pasampromptadaptersam] (49.9%). The results validate that our ELE-SAM also maintain the generalization capability. As illustrated in Fig. 9, ELE-SAM produces fewer artifacts and achieves more accurate segmentation, especially in complex background. It further underscore the robustness of ELE-SAM, making it a strong candidate for real-world applications which require high-quality segmentation without task-specific fine-tuning.
@@ 4.5 Ablation Study
In this section, we conduct ablation studies on ELE-40K and HQSeg-44K to investigate the effectiveness of High-Fidelity Mask Decoder (HFMD) and Context-Aware Prompt Adapter (CAPA). In HFMD, before the fusion process (the same as HQ-SAM [sam_hq]) between mask features and image fusion features, we plug in another light-weight refinement block (RB). The impacts of the additional refinement block and leveraging high-resolution mask features are discussed. Moreover, we showcase the effectiveness of CAPA over the original Prompt Adapter (PA).
@@
Impact of Different Backbones. When changing the size of ViT backbone, we find that ViT-L obtains the best performance on ELE-40K. The detailed metrics are listed in Tab. 7. While ViT-L outperforming ViT-B by a clear margin, the larger frozen ViT-H does not further promote the performance. It could be attributed to the over-smoothing issue [ru2023token] for deeper ViT.
Figure 11: Analysis on challenging cases. Example failure cases under highly challenging conditions (complex occlusions, low illumination, and adverse weather), which present opportunities for future robustness enhancement.
5 Limitation and Discussion
Although ELE-SAM achieves superior performance on the PTCHS task and general segmentation benchmarks, there is still potential for further improvement. As illustrated in Fig. 11, the segmentation quality may degrade under highly challenging conditions such as occlusion, low illumination, and adverse weather. While ELE-40K encompasses these scenarios, further architectural refinements can be developed to address these challenging cases.
In addition, as demonstrated in Tab. 6 and Tab. 7, ELE-SAM currently cannot achieve real-time segmentation based on large image encoders, such as ViT-L and ViT-H. Future work could leverage lightweight vision backbones to speed up inference and explore better multi-scale feature aggregation schemes to further promote the segmentation quality.
@@ 6 Conclusion
Acknowledgements
This work was financially supported by the State Grid Corporation Headquarters Science and Technology Project: Research on equipment operation and inspection disposal reasoning technology based on knowledge-enhanced generative model and intelligent agent and demonstration application (5700-202458333A-2-1-ZX). The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Data availability.
All datasets used in this paper are publicly available. ELE-40K can be accessed at https://github.com/Hhaizee/ELE-SAM. HQSeg-44K [sam_hq] can be accessed at https://github.com/SysCV/SAM-HQ. COCO [singh2024benchmarkingobjectdetectorscoco] can be accessed at https://github.com/cocodataset/cocoapi.